A particle of mass 100 g is thrown vertically
upwards with a speed of 5 m/s. The work done by
the force of gravity during the time the particle
goes up is-

Grade:10

2 Answers

kartikay guleria
11 years ago

-1.5 is the answer

renu akunuri
11 years ago

maximum height reached by the body which is thrown up vertically is h=(U/g)1/2

we know work=F.S

                        = mg.(u/g)1/2

                       = 0.1*(50)1/2

                        =0.5 * (2)1/2

                       =1.9 N.m